OpenStudy (explainitlikeimfive):

Graph the function in the interval from 0 to 2p.

OpenStudy (explainitlikeimfive):

\[y = \sin (\theta-2) \]

OpenStudy (anonymous):

y=sin(x-a) A shift of the sine graph a units to the right

OpenStudy (anonymous):

y=sin(x+a) A shift of the sine graph a units to the left

OpenStudy (anonymous):

y=sin(x)+a A shift of the sine grpah a units up y=sin(x)-a A shift of the sine graph a units down

OpenStudy (explainitlikeimfive):

From what I understand, I would pick C.

ganeshie8 (ganeshie8):

y=sin(x-2) shifts right by 2 units

ganeshie8 (ganeshie8):

strike off A and C cuz they have graphs shifted UP/DOWN

ganeshie8 (ganeshie8):

your answer is either B or D

OpenStudy (explainitlikeimfive):

I choose B.

ganeshie8 (ganeshie8):

thats correct ! but why, do u have a reason ? :)

OpenStudy (explainitlikeimfive):

Process of Elimination

ganeshie8 (ganeshie8):

why did u eliminate D

OpenStudy (anonymous):

Yep, these graphs look funny, the best way to determine it is to just plug in something like pi and see where it lands you on the grpah which is like sin(pi-2)=.9093 APPROX SO FIND ONE WHERE IT WAS

OpenStudy (explainitlikeimfive):

I answered D orginally, and it was wrong.

ganeshie8 (ganeshie8):

lol that surely is a good reason :)

OpenStudy (explainitlikeimfive):

Thanks to all.

OpenStudy (anonymous):

sorry my caps come on anmd off lol

OpenStudy (explainitlikeimfive):

I was wondering why you were yelling

ganeshie8 (ganeshie8):

as doulikepie says, just put \(\theta = 0\), then \(\sin (0 -2) = \sin (-2) \) = negative number since D has positive when \(\theta = 0\), we eliminate D

OpenStudy (explainitlikeimfive):

First come first served on medals. @doulikepiecauseidont

ganeshie8 (ganeshie8):

lol

OpenStudy (anonymous):

haha